# The consistency strength of determinacy when all sets are universally Baire

### v2

@@ Abstract.
The author gratefully acknowledges funding from L’ORÉAL Austria, in collaboration with the Austrian UNESCO Commission and in cooperation with the Austrian Academy of Sciences - Fellowship Determinacy and Large Cardinals. In addition, this research was funded in whole or in part by the Austrian Science Fund (FWF) [10.55776/V844, 10.55776/Y1498, 10.55776/I6087]. For open access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. Moreover, the author would like to thank Grigor Sargsyan for helpful discussions at the beginning of this project and the support through his NSF Career Award DMS-1352034 during a research visit at Rutgers University in the fall of 2019. Finally, the author is grateful for the very helpful comments and suggestions provided by the anonymous referee.
@@ 2. Preliminaries and notation
We start with the definition of universally Baire sets of reals, see for example [3, Definition 32.21] or [18, Definition 8.6].
@@
Analogous to the indexing in [13], the following extenders are indexed differently than in Mitchell-Steel indexing: Suppose $E$ is an extender on the sequence of a premouse $\mathcal{M}$ with critical point $\kappa$ that is a $j$-weak cutpoint and a limit of $j$-weak cutpoints for some $j<\omega$. Then $E$ is indexed at the successor of the least $j$-weak cutpoint above $\kappa$ of its own ultrapower, for the minimal such $j$. More precisely, a fine extender sequence in this article is defined as follows: We use the notation in [26]. In addition, if $E$ is an extender on the sequence of $\mathcal{M}=J_{\beta}^{\vec{E}}$ such that $\operatorname{crit}(E)$ is a $j$-weak cutpoint and a limit of $j$-weak cutpoints in $\mathcal{M}$ for some $j<\omega$, we say $E$ is of type S [^1] Here “S” stands for Sargsyan as this is a direct generalization of the indexing suggested by Sargsyan in [13]..
@@
A fine extender sequence in this article is a sequence $\vec{E}$ such that for each $\alpha\in\operatorname{dom}(\vec{E})$, $\vec{E}$ is acceptable at $\alpha$, and either $(\vec{E})_{\alpha}=\emptyset$, or $E_{\alpha}=(\vec{E})_{\alpha}$ is a $(\kappa,\alpha)$-pre-extender over $J_{\alpha}^{\vec{E}}$ for some $\kappa$ such that $J_{\alpha}^{\vec{E}}\models$$\kappa^{+}$ exists”, and
@@
    (Coherence) $i(\vec{E}\upharpoonright\kappa)\upharpoonright\alpha=\vec{E}\upharpoonright\alpha$ and $i(\vec{E}\upharpoonright\kappa)_{\alpha}=\emptyset$, where $i\colon J_{\alpha}^{\vec{E}}\rightarrow\operatorname{Ult}(J_{\alpha}^{\vec{E}},E_{\alpha})$ is the canonical embedding, and
@@ 3. Getting a translatable structure
In this section, we work in a model of “$\operatorname{AD}_{\mathbb{R}}+\text{ all sets of reals are uB}$” in order to define a hybrid structure that we are going to translate into a model with a cardinal $\lambda$ that is both a limit of Woodin cardinals and a limit of strong cardinals in the rest of this paper. Recall that by a result of Martin and Woodin, cf. [25, Theorem 9.1], $\operatorname{AD}_{\mathbb{R}}$ holds in every model of “AD+ all sets of reals are $\operatorname{AD}+\text{ all sets of reals are uB}$” as clearly every universally Baire set of reals is Suslin. The arguments in this section will come from descriptive inner model theory and will look familiar to readers who have seen $\operatorname{HOD}$ computations (cf., e.g. [29, 10, 12, 30]).
@@ 3.1. Hod premice
We consider hod premice as in [12, Definition 1.34] and just briefly recall some important notions here. Let $\mathcal{P}$ be a hod premouse. Then, as in [12] we let $\lambda^{\mathcal{P}}=o.t.(Y^{\mathcal{P}})$ and $(\delta_{\alpha}^{\mathcal{P}}\mid\alpha\leq\lambda^{\mathcal{P}})$ be the sequence of Woodin cardinals of $\mathcal{P}$ and their limits enumerated in increasing order. All $\delta_{\alpha}^{\mathcal{P}}$ will be strong cutpoints [^3] An ordinal $\gamma$ is called a strong cutpoint of a (hod) premouse $\mathcal{P}$ if there is no extender on the sequence of $\mathcal{P}$ such that $\operatorname{crit}(E)\leq\gamma<\operatorname{lh}(E)$. of $\mathcal{P}$ and our hod premice will always satisfy
@@
so they will for example look like [12, Figure 1.7.2] (an ordinary premouse, $\lambda^{\mathcal{P}}=0$), [12, Figure 1.7.3] ($\lambda^{\mathcal{P}}=1$), or [12, Figure 1.7.4] ($\lambda^{\mathcal{P}}=\omega$). As in [12, Definition 1.34], we use
@@
    $\mathcal{P}(i)$ to denote $\mathcal{P}$ up to its $i$’s level for $i\leq\lambda^{\mathcal{P}}$ and
@@
We will consider strategies for hod premice that satisfy hull condensation (cf., [12, Definition 1.31]).
@@
Let $\mathcal{P}$ be a hod premouse and let $\Sigma$ be an $(\omega_{1},\omega_{1}+1)$-iteration strategy for $\mathcal{P}$. Then we say $\Sigma$ has hull condensation if for any two stacks $\vec{\mathcal{T}}$ and $\vec{\mathcal{U}}$ on $\mathcal{P}$ such that $\vec{\mathcal{T}}$ is according to $\Sigma$ and $(\mathcal{P},\vec{\mathcal{U}})$ is a hull of $(\mathcal{P},\vec{\mathcal{T}})$ in the sense of [12, Definition 1.30], $\vec{\mathcal{U}}$ is according to $\Sigma$ as well.
Recall the definition of a hod pair (cf., [12, Definition 1.36]).
@@
    $(\mathcal{P},\Sigma)$ is an lsm pair (cf., [12, Definition 1.23]),
@@
As in [12, Definition 2.26], except that we require full iterability, we define the lower part closure relative to $\Sigma$ as follows. Note that this is denoted by the $\mathcal{W}$-operator in [13] and other articles.
@@
We define fullness preservation analogous to [12, Definition 2.27].
@@
Following [12, Section 4.3], we can now define direct limits of hod pairs as follows. For a hod pair $(\mathcal{P},\Sigma)$ such that $\Sigma$ has branch condensation (cf., [12, Definition 2.14]) and is fullness preserving, let
@@ 3.2. Suitable premice
We define suitable premice and related concepts as in [12, Section 4.1] but we will recall most notions for the reader’s convenience. The reader can consult [15, 30] for similar definitions.
@@
We define $(\kappa_{i},\Sigma)$-short tree, $(\kappa_{i},\Sigma)$-maximal tree and $(\kappa_{i},\Sigma)$-correctly guided finite (mixed) stack , the last model of a $(\kappa_{i},\Sigma)$-correctly guided finite mixed stack, $S(\kappa_{i},\Sigma)$, $F(\kappa_{i},\Sigma)$, and $f$-iterability as in [12, Definitions 4.3 - 4.7], adapted to our definition of suitability. We recall the concepts here for the reader’s convenience. In what follows, let $\mathcal{P}$ be a hod premouse that is $\Sigma$-suitable at some $\kappa_{i}$, $i<\omega$, for some fullness preserving iteration strategy $\Sigma$. Write
@@
As usual, $f$-iterability embeddings for a given $\vec{\mathcal{T}}$ are independent of the choice of $\vec{b}$ (see [12, Lemma 4.8]). Moreover, we can define a strengthening, called strong $f$-iterability, where this uniqueness also does not depend on the choice of the stack $\vec{\mathcal{T}}$ (see [12, Definition 4.10]).
In the previous definitions we omit the reference to $\kappa_{i}$ and $\Sigma$ if $i=0$ and $\Sigma=\emptyset$ or it is clear from the context to which $i<\omega$ and strategy $\Sigma$ we are referring. As in [12, Section 4.2] we can define when a $(\kappa_{i},\Sigma)$-suitable premouse is $A$-iterable [^6] Sargsyan calls this $B$-iterability. for an $\operatorname{OD}_{\Sigma}$ set of reals $A$:
@@ 3.3. Properties of iteration strategies
We will need the following standard properties of iteration strategies. The first one is super fullness preservation. This notion goes back to [12, Definition 3.33] and is also used in [13, Definition 2.3]. We simply adapt the definition to our context. Recall that we will specify the models in which we apply the definitions from this section later. In this section we sometimes explicitly mention a model $M$ when we discuss generic closure of this model under a strategy and related notions. This will later be applied to models $M$ that are strategy mice.
@@
    for every iteration $j\colon\mathcal{P}\rightarrow\mathcal{Q}$ according to $\Sigma\upharpoonright V_{\eta}$, every self well-ordered set $X\in V_{\eta}$, and every $\operatorname{Col}(\omega,j(\delta^{n}))$-generic $g$ over $\mathcal{Q}$ such that $X\in\operatorname{HC}^{\mathcal{Q}[g]}$, letting $x\in\mathbb{R}^{\mathcal{Q}[g]}$ be generic over $L_{\omega}[X]$ and coding $X$, [^7] For the definition of $\mathcal{P}$-constructions for hod premice see $\mathcal{S}$-constructions in [12, Section 3.8]. As such constructions are due to Steel, they are called $\mathcal{S}$-constructions in [12]. We nevertheless decided to use the more common terminology $\mathcal{P}$-construction here.
@@
Now, following [13, Definition 2.4], we can define that a model $M$ is generically closed under $\Sigma$ if $\Sigma$ is super fullness preserving in all generic extensions of $M$. Formally, this is defined as follows.
@@
Now, the following definition adapts [13, Definition 2.6] to our context (cf., [12, Definition 4.14]).
@@ 3.4. Translatable structures
Suppose first that $\Theta>\theta_{\omega}$. Then, by [12, Lemma 0.23], there is a hod pair $(\mathcal{P},\Sigma_{\mathcal{P}})$ such that $\Sigma_{\mathcal{P}}$ has branch condensation, [^9] For branch condensation in the sense of [12, Definition 2.14] it is essential that we are working with hod mice under the minimality assumption of [12]. This suffices for our aim but we recommend that the reader interested in a more general setting consults [27]. is fullness-preserving and $V_{\theta_{\omega}}^{\operatorname{HOD}}$ is a shortening of the direct limit of all $\Sigma_{\mathcal{P}}$-iterates of $\mathcal{P}$, i.e., $\mathcal{M}_{\infty}(\mathcal{P},\Sigma_{\mathcal{P}})|\theta_{\omega}=V_{\theta_{\omega}}^{\operatorname{HOD}}.$ In this case, we let
@@
Here $w(Code(\Sigma_{\mathcal{P}}))$ denotes the Wadge-rank of the canonical code of $\Sigma_{\mathcal{P}}$ as a set of reals. Note that $\mathcal{M}_{\infty}$ is well-defined as the direct limit models $\mathcal{M}_{\infty}(\mathcal{P},\Sigma_{\mathcal{P}})$ line up and do not depend on the choice of $(\mathcal{P},\Sigma_{\mathcal{P}})$ (cf. [12, Lemma 0.23 and Theorem 4.24]). In fact, by the results of [12], $V_{\Theta}^{\operatorname{HOD}}=\mathcal{M}_{\infty}$ and there is a canonical iteration strategy, which we call $\Sigma$, for $\mathcal{M}_{\infty}$ given by the join of the strategies for its hod initial segments (see also [30, Section 3] on how $\mathcal{M}_{\infty}$ can be represented as a direct limit of a directed system of hod pairs).
@@
Universal Baireness ensures that we can canonically extend each $\Sigma_{i}$ to an iteration strategy acting on stacks of normal iteration trees of length $\lambda$ for any ordinal $\lambda$. See [11, Lemma 1.21] for this argument and in particular for how iteration strategies need to be coded as sets of reals to make the coding sufficiently absolute. We also write $\Sigma_{i}$ for this canonical extension of the original strategy $\Sigma_{i}$ of $\mathcal{M}_{\infty}$ between $\delta_{i-1}$ and $\delta_{i}$ as well as $\Sigma$ for the resulting extension of the original strategy $\Sigma$ of $\mathcal{M}_{\infty}$.
@@
Moreover, by our choice of the strategies $\Sigma_{i}$ (in particular, by the fact that they have branch condensation in the sense of [12, Definition 2.14]) and generic interpretability for hod mice (see [12, Theorem 3.10]) we have the following lemma.
@@
Let $\mathcal{W}$ be a proper class hybrid strategy premouse in the sense of [12, Chapter 1]. Then $\mathcal{W}$ is a translatable structure iff there is a sequence of ordinals $(\delta_{i}\mid i<\omega)$ such that $\mathcal{W}$ satisfies the following conditions.
@@
    $\mathcal{W}\vDash$ “there is no inner model with a superstrong cardinal”.
@@
    Let $\delta_{\omega}$ be the limit of $\delta_{i}$, $i<\omega$. Let $G$ be $\operatorname{Col}(\omega,{<}\delta_{\omega})$-generic over $\mathcal{W}$ and let $M$ be the derived model computed in $\mathcal{W}(\mathbb{R}^{*})$. Let $\Phi=(\Sigma^{\mathcal{W}})^{G}\upharpoonright\operatorname{HC}^{M}$ for $\Sigma^{\mathcal{W}}=\bigoplus_{i<\omega}\Sigma_{i}$. Then $\Phi\in M$ and in $M$, $\mathcal{W}|(\delta_{i}^{+\omega})^{\mathcal{W}}$ is a $\kappa_{i}$-suitable $\emptyset$-iterable $\Phi$-premouse (cf. [13, Definition 2.7] for the definition of $\emptyset$-iterability).
@@
Suppose otherwise, i.e., there is a level $J_{\xi}^{\Sigma}[\mathcal{M}_{\infty}]$ that projects across $\mathcal{M}_{\infty}\cap\operatorname{Ord}$. That means, $J_{\xi}^{\Sigma}[\mathcal{M}_{\infty}]$ projects across some $\delta_{i+1}$, for $i<\omega$. But then $J_{\xi+1}^{\Sigma}[\mathcal{M}_{\infty}]$ shows up in the direct limit system giving rise to $\mathcal{M}_{\infty}$, contradicting the fact that there is a new subset of $\mathcal{M}_{\infty}$ in $J_{\xi+1}^{\Sigma}[\mathcal{M}_{\infty}]$. ∎

### v3

@@ Abstract.
The author gratefully acknowledges funding from L’ORÉAL Austria, in collaboration with the Austrian UNESCO Commission and in cooperation with the Austrian Academy of Sciences - Fellowship Determinacy and Large Cardinals. In addition, this research was funded in part by the Austrian Science Fund (FWF) [10.55776/V844, 10.55776/Y1498, 10.55776/I6087]. For open access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. Moreover, the author would like to thank Grigor Sargsyan for helpful discussions at the beginning of this project and the support through his NSF Career Award DMS-1352034 during a research visit at Rutgers University in the fall of 2019. In addition, the author would like to thank Farmer Schlutzenberg, Benjamin Siskind, and Lena Wallner for helpful comments on an earlier version of this article. Finally, the author is grateful for the very helpful comments and suggestions provided by the anonymous referee.
@@ 1. Introduction
As a corollary to our proof we obtain a proof of Steel’s unpublished result [24] on the exact consistency strength lower bound for $\operatorname{AD}_{\mathbb{R}}$ mentioned at the beginning of this introduction. We thank John Steel for his permission to include it here.
Corollary 1.4 (Steel).
Suppose there is a proper class model of $\operatorname{AD}_{\mathbb{R}}$. Then the $\operatorname{AD}_{\mathbb{R}}$-hypothesis is consistent, i.e., there is a transitive model $\mathcal{M}$ of $\operatorname{ZFC}$ containing all ordinals such that $\mathcal{M}$ has a cardinal $\lambda$ that is a limit of Woodin cardinals and a limit of ${<}\lambda$-strong cardinals.
The corollary follows from our argument by observing that Sargsyan’s [11, 13] yields a hod mouse with a limit of Woodin cardinals $\lambda$ and strategies up to $\lambda$ from $\operatorname{AD}_{\mathbb{R}}$. Restricting the translation procedure of this article to strategies below $\lambda$ results in a transitive model $\mathcal{M}$ of $\operatorname{ZFC}$ containing all ordinals such that $\mathcal{M}$ has a cardinal $\lambda$ that is a limit of Woodin cardinals and a limit of ${<}\lambda$-strong cardinals. Steel’s original unpublished proof was different from ours. He, assuming $\operatorname{AD}_{\mathbb{R}}$, Prikry-forced a hod mouse (in the sense of [11]) $H$ whose derived model in $V$ and then translated it into a model of the $\operatorname{AD}_{\mathbb{R}}$-hypothesis whose derived model is still $V$. His proof that the translation works used that certain mouse operators are equivalent on a cone, and that the bases of those cones have been made generic over $H$ at the appropriate places.
@@ 2. Preliminaries and notation
We start with the definition of universally Baire sets of reals; see, for example, [3, Definition 32.21] or [19, Definition 8.6].
@@
Analogous to the indexing in [14], the following extenders are indexed differently than in Mitchell-Steel indexing: Suppose $E$ is an extender on the sequence of a premouse $\mathcal{M}$ with critical point $\kappa$ that is a $j$-weak cutpoint and a limit of $j$-weak cutpoints for some $j<\omega$. Then $E$ is indexed at the successor of the least $j$-weak cutpoint above $\kappa$ of its own ultrapower, for the minimal such $j$. More precisely, a fine extender sequence in this article is defined as follows: We use the notation in [28]. In addition, if $E$ is an extender on the sequence of $\mathcal{M}=J_{\beta}^{\vec{E}}$ such that $\operatorname{crit}(E)$ is a $j$-weak cutpoint and a limit of $j$-weak cutpoints in $\mathcal{M}$ for some $j<\omega$, we say $E$ is of type S [^1] Here “S” stands for Sargsyan as this is a direct generalization of the indexing suggested by Sargsyan in [14].. Another difference is that, similar as in [22], we do not require coherence above $\lambda$, the limit of Woodin and strong cardinals we aim to obtain. The reason for this is that in the model we construct there will be extenders indexed throughout the ordinals but no extenders on the sequence have critical point $\geq\lambda$. So coherence cannot hold above $\lambda$.
@@
For an ordinal $\lambda$, a fine extender sequence up to $\lambda$ in this article is a sequence $\vec{E}$ such that for each $\alpha\in\operatorname{dom}(\vec{E})$, $\vec{E}$ is acceptable at $\alpha$, and either $(\vec{E})_{\alpha}=\emptyset$, or $E_{\alpha}=(\vec{E})_{\alpha}$ is a $(\kappa,\alpha)$-pre-extender over $J_{\alpha}^{\vec{E}}$ for some $\kappa$ such that $J_{\alpha}^{\vec{E}}\models$$\kappa^{+}$ exists”, and
@@
    (Coherence) for $\alpha<\lambda$, $i(\vec{E}\upharpoonright\kappa)\upharpoonright\alpha=\vec{E}\upharpoonright\alpha$ and $i(\vec{E}\upharpoonright\kappa)_{\alpha}=\emptyset$, where $i\colon J_{\alpha}^{\vec{E}}\rightarrow\operatorname{Ult}(J_{\alpha}^{\vec{E}},E_{\alpha})$ is the canonical embedding, and
@@ 3. Getting a translatable structure
In this section, we work in a model of “$\operatorname{AD}_{\mathbb{R}}+\text{ all sets of reals are uB}$” in order to define a hybrid structure that we are going to translate into a model with a cardinal $\lambda$ that is both a limit of Woodin cardinals and a limit of strong cardinals in the rest of this paper. Recall that by a result of Martin and Woodin, cf. [27, Theorem 9.1], $\operatorname{AD}_{\mathbb{R}}$ holds in every model of “AD+ all sets of reals are $\operatorname{AD}+\text{ all sets of reals are uB}$” as clearly every universally Baire set of reals is Suslin. The arguments in this section will come from descriptive inner model theory and will look familiar to readers who have seen $\operatorname{HOD}$ computations (cf., e.g. [31, 10, 13, 32]).
@@ 3.1. Hod premice
We consider hod premice as in [13, Definition 1.34] and just briefly recall some important notions here. Let $\mathcal{P}$ be a hod premouse. Then, as in [13] we let $\lambda^{\mathcal{P}}$ be the order type of the Woodin cardinals of $\mathcal{P}$ and their limits [^3] In the notation of [13], $\lambda^{\mathcal{P}}=o.t.(Y^{\mathcal{P}})$. and $(\delta_{\alpha}^{\mathcal{P}}\mid\alpha\leq\lambda^{\mathcal{P}})$ be the sequence of Woodin cardinals of $\mathcal{P}$ and their limits enumerated in increasing order. All $\delta_{\alpha}^{\mathcal{P}}$ will be strong cutpoints [^4] An ordinal $\gamma$ is called a strong cutpoint of a (hod) premouse $\mathcal{P}$ if there is no extender on the sequence of $\mathcal{P}$ such that $\operatorname{crit}(E)\leq\gamma<\operatorname{lh}(E)$. of $\mathcal{P}$ and our hod premice will always satisfy
@@
so they will for example look like [13, Figure 1.7.2] (an ordinary premouse, $\lambda^{\mathcal{P}}=0$), [13, Figure 1.7.3] ($\lambda^{\mathcal{P}}=1$), or [13, Figure 1.7.4] ($\lambda^{\mathcal{P}}=\omega$). As in [13, Definition 1.34], we use
@@
    $\mathcal{P}(i)$ to denote $\mathcal{P}$ up to its $i$’s level [^5] So, in our situation, $\mathcal{P}(i)=\mathcal{P}|(\delta_{i}^{\mathcal{P}})^{+\omega}$. for $i\leq\lambda^{\mathcal{P}}$ and
@@
We will consider strategies for hod premice that satisfy hull condensation (cf., [13, Definition 1.31]).
@@
Let $\mathcal{P}$ be a hod premouse and let $\Sigma$ be an $(\omega_{1},\omega_{1}+1)$-iteration strategy for $\mathcal{P}$. Then we say $\Sigma$ has hull condensation if for any two stacks $\vec{\mathcal{T}}$ and $\vec{\mathcal{U}}$ on $\mathcal{P}$ such that $\vec{\mathcal{T}}$ is according to $\Sigma$ and $(\mathcal{P},\vec{\mathcal{U}})$ is a hull of $(\mathcal{P},\vec{\mathcal{T}})$ in the sense of [13, Definition 1.30], $\vec{\mathcal{U}}$ is according to $\Sigma$ as well.
Recall the definition of a hod pair (cf., [13, Definition 1.36]).
@@
    $(\mathcal{P},\Sigma)$ is an lsm pair (cf., [13, Definition 1.23]),
@@
As in [13, Definition 2.26], except that we require full iterability, we define the lower part closure relative to $\Sigma$ as follows. Note that this is denoted by the $\mathcal{W}$-operator in [14] and other articles.
@@
We define fullness preservation analogous to [13, Definition 2.27].
@@
Following [13, Section 4.3], we can now define direct limits of hod pairs as follows. For a hod pair $(\mathcal{P},\Sigma)$ such that $\Sigma$ has branch condensation (cf., [13, Definition 2.14]) and is fullness preserving, let
@@ 3.2. Suitable premice
We define suitable premice and related concepts as in [13, Section 4.1] but we will recall most notions for the reader’s convenience. The reader can consult [16, 32] for similar definitions.
@@
We define $(\kappa_{i},\Sigma)$-short tree, $(\kappa_{i},\Sigma)$-maximal tree and $(\kappa_{i},\Sigma)$-correctly guided finite (mixed) stack , the last model of a $(\kappa_{i},\Sigma)$-correctly guided finite mixed stack, $S(\kappa_{i},\Sigma)$, $F(\kappa_{i},\Sigma)$, and $f$-iterability as in [13, Definitions 4.3 - 4.7], adapted to our definition of suitability. We recall the concepts here for the reader’s convenience. In what follows, let $\mathcal{P}$ be a hod premouse that is $\Sigma$-suitable at some $\kappa_{i}$, $i<\omega$, for some fullness preserving iteration strategy $\Sigma$. Write
@@
As usual, $f$-iterability embeddings for a given $\vec{\mathcal{T}}$ are independent of the choice of $\vec{b}$ (see [13, Lemma 4.8]). Moreover, we can define a strengthening, called strong $f$-iterability, where this uniqueness also does not depend on the choice of the stack $\vec{\mathcal{T}}$ (see [13, Definition 4.10]).
In the previous definitions we omit the reference to $\kappa_{i}$ and $\Sigma$ if $i=0$ and $\Sigma=\emptyset$ or it is clear from the context to which $i<\omega$ and strategy $\Sigma$ we are referring. As in [13, Section 4.2] we can define when a $(\kappa_{i},\Sigma)$-suitable premouse is $A$-iterable [^8] Sargsyan calls this $B$-iterability. for an $\operatorname{OD}_{\Sigma}$ set of reals $A$:
@@ 3.3. Properties of iteration strategies
We will need the following standard properties of iteration strategies. The first one is super fullness preservation. This notion goes back to [13, Definition 3.33] and is also used in [14, Definition 2.3]. We simply adapt the definition to our context. Recall that we will specify the models in which we apply the definitions from this section later. In this section we sometimes explicitly mention a model $M$ when we discuss generic closure of this model under a strategy and related notions. This will later be applied to models $M$ that are strategy mice.
@@
    for every iteration $j\colon\mathcal{P}\rightarrow\mathcal{Q}$ according to $\Sigma\upharpoonright V_{\eta}$, every self well-ordered set $X\in V_{\eta}$, and every $\operatorname{Col}(\omega,j(\delta^{n}))$-generic $g$ over $\mathcal{Q}$ such that $X\in\operatorname{HC}^{\mathcal{Q}[g]}$, letting $x\in\mathbb{R}^{\mathcal{Q}[g]}$ be generic over $L_{\omega}[X]$ and coding $X$, [^9] For the definition of $\mathcal{P}$-constructions for hod premice see $\mathcal{S}$-constructions in [13, Section 3.8]. As such constructions are due to Steel, they are called $\mathcal{S}$-constructions in [13]. We nevertheless decided to use the more common terminology $\mathcal{P}$-construction here.
@@
Now, following [14, Definition 2.4], we can define that a model $M$ is generically closed under $\Sigma$ if $\Sigma$ is super fullness preserving in all generic extensions of $M$. Formally, this is defined as follows.
@@
Now, the following definition adapts [14, Definition 2.6] to our context (cf., [13, Definition 4.14]).
@@ 3.4. Translatable structures
Suppose first that $\Theta>\theta_{\omega}$. Then, by [13, Lemma 0.23], there is a hod pair $(\mathcal{P},\Sigma_{\mathcal{P}})$ such that $\Sigma_{\mathcal{P}}$ has branch condensation, [^11] For branch condensation in the sense of [13, Definition 2.14] it is essential that we are working with hod mice under the minimality assumption of [13]. This suffices for our aim but we recommend that the reader interested in a more general setting consults [29]. is fullness-preserving and $V_{\theta_{\omega}}^{\operatorname{HOD}}$ is a shortening of the direct limit of all $\Sigma_{\mathcal{P}}$-iterates of $\mathcal{P}$, i.e., $\mathcal{M}_{\infty}(\mathcal{P},\Sigma_{\mathcal{P}})|\theta_{\omega}=V_{\theta_{\omega}}^{\operatorname{HOD}}.$ In this case, we let
@@
Here $w(Code(\Sigma_{\mathcal{P}}))$ denotes the Wadge-rank of the canonical code of $\Sigma_{\mathcal{P}}$ as a set of reals. Note that $\mathcal{M}_{\infty}$ is well-defined as the direct limit models $\mathcal{M}_{\infty}(\mathcal{P},\Sigma_{\mathcal{P}})$ line up and do not depend on the choice of $(\mathcal{P},\Sigma_{\mathcal{P}})$ (cf. [13, Lemma 0.23 and Theorem 4.24]). In fact, by the results of [13], $V_{\Theta}^{\operatorname{HOD}}=\mathcal{M}_{\infty}$ and there is a canonical iteration strategy, which we call $\Sigma$, for $\mathcal{M}_{\infty}$ given by the join of the strategies for its hod initial segments (see also [32, Section 3] on how $\mathcal{M}_{\infty}$ can be represented as a direct limit of a directed system of hod pairs).
@@
Universal Baireness ensures that we can canonically extend each $\Sigma_{i}$ to an iteration strategy acting on stacks of normal iteration trees of length $\lambda$ for any ordinal $\lambda$. See [12, Lemma 1.21] for this argument and in particular for how iteration strategies need to be coded as sets of reals to make the coding sufficiently absolute. We also write $\Sigma_{i}$ for this canonical extension of the original strategy $\Sigma_{i}$ of $\mathcal{M}_{\infty}$ between $\delta_{i-1}$ and $\delta_{i}$ as well as $\Sigma$ for the resulting extension of the original strategy $\Sigma$ of $\mathcal{M}_{\infty}$.
@@
Moreover, by our choice of the strategies $\Sigma_{i}$ (in particular, by the fact that they have branch condensation in the sense of [13, Definition 2.14]) and generic interpretability for hod mice (see [13, Theorem 3.10]) we have the following lemma.
@@
Let $\mathcal{W}$ be a proper class hybrid strategy premouse in the sense of [13, Chapter 1]. Then $\mathcal{W}$ is a translatable structure iff there is a sequence of ordinals $(\delta_{i}\mid i<\omega)$ such that $\mathcal{W}$ satisfies the following conditions.
@@
    $\mathcal{W}\vDash$ “there is no inner model with a superstrong cardinal”. [^12] In this article, we are aiming for the consistency of a theory weaker than the existence of a superstrong cardinal, so we will never consider any inner models not satisfying (5) and this condition is redundant here.
@@
    Let $\delta_{\omega}$ be the limit of $\delta_{i}$, $i<\omega$. Let $G$ be $\operatorname{Col}(\omega,{<}\delta_{\omega})$-generic over $\mathcal{W}$ and let $M$ be the derived model computed in $\mathcal{W}(\mathbb{R}^{*})$. Let $\Phi=(\Sigma^{\mathcal{W}})^{G}\upharpoonright\operatorname{HC}^{M}$ for $\Sigma^{\mathcal{W}}=\bigoplus_{i<\omega}\Sigma_{i}$. Then $\Phi\in M$ and in $M$, $Lp_{\omega}^{\infty}(\mathcal{W}|\delta_{i})$ is a $\kappa_{i}$-suitable $\emptyset$-iterable $\Phi$-premouse (cf. [14, Definition 2.7] for the definition of $\emptyset$-iterability).
@@
Suppose otherwise, i.e., there is a $\xi\geq 1$ such that $J_{\xi}^{\Sigma}[\mathcal{M}_{\infty}]$ projects across $\mathcal{M}_{\infty}\cap\operatorname{Ord}$. Then $J_{\xi}^{\Sigma}[\mathcal{M}_{\infty}]$ is an anomalous hod premouse of type III (see Definition 3.23 in [13]) and using fullness preservation of $\Sigma$ we can follow the proof of Theorem 6.1 in [13] (see also the argument for Claim 2 in the proof of Lemma 6.23 in [13]) to get a contradiction. ∎